OpenStudy (pokemon23):

Express in simplest form the area of a trapezoid whose bases are 11 and 28 and whose height is represented is represented by 4y. Express in simplest form the area of a trapezoid whose bases are 11 and 28 and whose height is represented is represented by 4y. @Mathematics

OpenStudy (pokemon23):

TranceNova how do you look for how many exp you have?

OpenStudy (pokemon23):

thanks maiden for helping me

OpenStudy (anonymous):

 What we have in the picture is 2 triangles and a rectangle. We know the area of a triangle is 1/2*b*h and the area of a rectangle is l*h. The length of the rectangle is 11. The base of both triangles can be solved using this equation: 2x + 11 = 28, where x is the base of the triangle. 2x = 17 x = 17/2 So the area of the trapezoid is 2*1/2b*h + l*h A = 17*4y/2 + 11*4y = 34y + 44y = 78y A = 78y